The cost in dollars of making x items is given by the function


C(x) = 10x + 500.


a. The fixed cost is determined when zero items are produced. Find the fixed cost for this item.


b. What is the cost of making 25 items?


c. Suppose the maximum cost allowed is $1500. What are the domain and range of the cost function,


C(x)?



1
Expert's answer
2021-12-02T12:18:17-0500

a) Fixed cost

= C(x) = 10 x 0 + 500

C (x) = 500$


b) Cost for 25 items


C(x) = 10x 25 + 500


= 250 + 500

Answer= 750$


d) Cost function.

1500$ = 10x + 500

1500 -500 = 10x

1000 = 10x

Thus x =100

Answer= In order to attain a maximum cost of 1500$ the domain range should be between 0 to 100 Items.

thus 0<= x >= 100


The minimum cost is 500$ and the maximum cost is 1500$


thus 500$ <= C(x) >= 1500$